OpenStudy (anonymous):

Only one question I cannot figue out. can someone help please. What are the solutions? 1/2x^2+ 2x + 3 =0

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

subtract 3 from both sides

OpenStudy (anonymous):

what like terms, sorry I am having a lot of problems with this.

OpenStudy (anonymous):

lol dude its ok, equations can be confusing to some

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x^2+2x+3=0 thats what ur starting out with\]

OpenStudy (anonymous):

thats what ur starting with *

OpenStudy (anonymous):

right. Than you want to get the three to the other side right.

OpenStudy (anonymous):

u can do that first. the thing with equations is that there is no direct way of solving it. there r many ways. just make sure what ever u are doing, make sure it mathimatically makes sense

OpenStudy (anonymous):

1/2^2+2x=-3

OpenStudy (anonymous):

good, now simplify 1/2^x+2x

OpenStudy (anonymous):

here is my problem and were I am stuck. Should I try and get rid of the 1/2?

OpenStudy (anonymous):

do u have a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

square the half

OpenStudy (anonymous):

lost me again. Sorry.

OpenStudy (anonymous):

lol its ok. do u have a graphing calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

scientific?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. put 1 divided by 2 and square it

OpenStudy (anonymous):

it should be 1/2^2

OpenStudy (anonymous):

there should be a calculator online

OpenStudy (anonymous):

0.25

OpenStudy (anonymous):

which is 1/4

OpenStudy (anonymous):

I really need help on this to finish a test! last question.

OpenStudy (anonymous):

@freak098 , you clearly do not understand the question - none of what you said is valid here.

OpenStudy (anonymous):

This is a quadratic equation, and doesn't look factorable at first glance. Quadratic formula is your best option.

OpenStudy (anonymous):

I think you use the quadratic formula. Am I correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Quadratic formula always works, so is typically the weapon of choice in this situation. A good way to start though, is to multiply everything by 2 first to clear the fraction.

OpenStudy (anonymous):

2(1/2x^2+2x+3=0(2) x^2+4x+6=0 correct?

OpenStudy (anonymous):

Yeah, then use the QF on that.

OpenStudy (anonymous):

ok. I think I got it now. It was the 1/2 I was confused on. Thank you very much.

OpenStudy (anonymous):

No problem. When I tried it, I was getting non-real answers, so don't be surprised about that.

OpenStudy (anonymous):

ok. I will try it